xx WaveToyF77

WaveToyF77

Abstract

1 Introduction 2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References

6 Parameters

11bound 10Scope: restrictedx01  KEYWORD

11Description: Type of boundary condition to use

11 Range 10 x01 Default: none
11none  10Apply no boundary condition
11flat  10Flat (von Neumann, n grad phi = 0) boundary condition
11static 10Static (Dirichlet, dphi/dt=0) boundary condition
11radiation 10Radiation boundary condition
11robin  10Robin (phi(r) = C/r) boundary condition
11zero  10Zero (Dirichlet, phi=0) boundary condition

11

7 Interfaces

General Implements: wavetoy

Inherits: grid

Grid Variables

7.0.1 PUBLIC GROUPS

11x Group Names x10x Variable Names xx01 Details x11x
11 scalarevolve 10  x01  compact110
11                 10  phi  x01 description11The evolved scalar field
11                 10  x01 dimensions113
11                 10  x01 distribution11DEFAULT
11                 10  x01 group type11GF
11                 10  x01 timelevels113
11                 10  x01variable type11REAL

11

Uses header: Symmetry.h 8 Schedule This section lists all the variables which are assigned storage by thorn CactusWave/WaveToyF77. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11Always:  11x
11scalarevolve[3]11x

11x  11x

11

Scheduled Functions

CCTKxSTARTUP x wavetoyf77xstartup x register banner x

11x11Language:11fortran
11 x 11Type: 11function

11

CCTKxBASEGRID x wavetoyf77xinitsymbound x schedule symmetries x

11x11Language:11fortran
11 x 11Type: 11function

11

CCTKxPOSTxRECOVERxVARIABLES x applybcs x apply boundary conditions x

11x11After:11wavetoyxboundaries
11 x 11Type:11group

11

CCTKxEVOL x wavetoyf77xevolution x evolution of 3d wave equation x

11x11Language:11fortran
11 x 11Type: 11function

11

CCTKxEVOL x wavetoyf77xboundaries x boundaries of 3d wave equation x

11x11After: 11wavetoyxevolution
11 x 11Language:11fortran
11 x 11Options: 11level
11 x 11Sync:  11scalarevolve
11 x 11Type: 11function

11

CCTKxEVOL x applybcs x apply boundary conditions x

11x11After:11wavetoyxboundaries
11 x 11Type:11group

11

CCTKxPOSTRESTRICT x wavetoyf77xboundaries x boundaries of 3d wave equation x

11x11Language:11fortran
11 x 11Options: 11level
11 x 11Sync:  11scalarevolve
11 x 11Type: 11function

11

CCTKxPOSTRESTRICT x applybcs x apply boundary conditions x

11x11After:11wavetoyxboundaries
11 x 11Type:11group

11

CCTKxPOSTRESTRICTINITIAL x wavetoyf77xboundaries x boundaries of 3d wave equation

11x11Language:11fortran
11 x 11Options: 11level
11 x 11Sync:  11scalarevolve
11 x 11Type: 11function

11

CCTKxPOSTRESTRICTINITIAL x applybcs x apply boundary conditions x

11x11After:11wavetoyxboundaries
11 x 11Type:11group

11

CCTKxPOSTxRECOVERxVARIABLES x wavetoyf77xboundaries x boundaries of 3d wave equation x

11x11Language:11fortran
11 x 11Options: 11level
11 x 11Sync:  11scalarevolve
11 x 11Type: 11function

11

Aliased Functions x

11Alias Name: xxxxxxx 11Function Name:
11ApplyBCs  11WaveToyF77xApplyBCs
11WaveToyF77xBoundaries11WaveToyxBoundaries
11WaveToyF77xEvolution 11WaveToyxEvolution